3x-5(3x-7)=2(x+9)+45
Answer by ankor@dixie-net.com(22740)

3x - 5(3x-7) = 2(x+9)+ 45
:
Mult what's inside the brackets, remember if you have a negative outside the
brackets you change the signs inside the brackets:
3x - 15x + 35 = 2x + 18 + 45
:
Perform the operations necessary to get the variables (x) on the left and the numbers on the right
Subtract 35 from both sides:
3x - 15x = 2x + 18 + 45 - 35
:
Subtract 2x from both sides, do the math on the right:
3x - 15x - 2x = 28
:
Do the math on the x's
-14x = 28
:
We want x to be positive, mult eq by -1, changes all the signs
+14x = -28
:
Divide both sides by 14 and you have:
x = -28/14
x = -2
:
:
Check our solution in the original equation, substitute -2 for x:
3x - 5(3x-7) = 2(x+9) + 45
:
3(-2) - 5(3(-2)-7) = 2((-2)+9) + 45
-6 - 5(-6-7) = 2(7) + 45
-6 - 5(-13) = 14 + 45
-6 + 65 = 59
59 = 59 confirms our solution